Most graduate schools of business require applicants for admission to take the Graduate Management Admission Council’s GMAT examination. Scores on the GMAT are roughly normally distributed with a mean of 506 and a standard deviation of 96. (14 pts)


a. What is the probability of an individual scoring above 520? (with illustration, 6 pts)


b. What is the probability of an individual scoring below 506? (with illustration, 4 pts)


c. What is the probability of an individual scoring from 387 to 712? (with illustration, 4 pts)


1
Expert's answer
2022-05-04T12:59:52-0400

a.


"P(X>520)=1-P(Z\le\dfrac{520-506}{96})\approx0.4420"

b.


"P(X<506)=P(Z<\dfrac{506-506}{96})"




"=P(Z<0)=0.5"

c.


"P(387<X<712)=P(Z<\dfrac{712-506}{96})"




"-P(Z\le\dfrac{387-506}{96})\approx0.8765"